OpenStudy (anonymous):

an investor invested a total of $1700 in two mutual funds. one fund earned a 5% profit value while the other earned 2% profit. if the investor's profit was $67, how much was in each mutual fund? the amount invested in the fund that earned 5% was __________. the amount invested in the fund that earned 2% was ________

OpenStudy (ankit042):

let us assume investment in the first be x so total profit 0.05x + 0.02(1700-x) = 67 Now solve for x

OpenStudy (anonymous):

85

OpenStudy (anonymous):

i think

OpenStudy (ankit042):

are you sure?

OpenStudy (ankit042):

try solving the equation

OpenStudy (anonymous):

i dont get it

OpenStudy (ranga):

Solve the equation mentioned by ankit above: 0.05x + 0.02(1700-x) = 67 0.05x + (0.02)(1700) - 0.02x = 67 0.03x + 34 = 67 0.03x = 33 x = 1100 So the amount invested in the fund that earned 5% was $1100. The amount invested in the fund that earned 2% was $600.